OpenStudy (anonymous):

Find two integers that make the equation true. x2 = 16 a. 4, -4 c. 14, 2 b. 2, 8 d. 7, 9

OpenStudy (anonymous):

come on man easy.A

OpenStudy (anonymous):

It's a because 4^2 = 4*4 = 16 and -4^2 = -4*-4 = 16.